Toms of Maine sells toothpaste in a tube advertised to contain 8 ounces. In their continuous production process, the amount of toothpaste put in a tube is normally distributed with a mean of 8.10 ounces and a standard deviation of 0.05 ounces. If the actual capacity of the tubes used is 8.22 ounces, what proportion of the tubes will be filled beyond capacity?
Answer #1

Process output, X ~ Normal(μ=8.10, σ=0.05)

We have to find the probability P(X > 8.22)

P(X > 8.22) = 1 - P(X < 8.22)

= 1 - P{(X - μ)/σ < (8.22 - μ)/σ}

= 1 - P(Z < (8.22 - 8.10)/0.05) Where Z ~ Normal (0,1)

= 1 - P(Z < 2.4)

= 1 - F(2.4)

Check the standard normal table to find the cumulative distribution of z = 2.4 i.e. F(2.4) = 0.99180

STANDARD NORMAL DISTRIBUTION: Table Values Represent AREA to the LEFT of the Z score Z 00 01 03 .04 05 06 07 .08 09 02 52392

So,

P(X > 8.22) = 1 - F(2.4) = 1 - 0.99180 = 0.0082